OpenStudy (anonymous):

In a polynomial function, is the constant always the y intercept? For example: f(x) = x^3 - 33x^2 +216x - 2 Is the y intercept at y=-2?

OpenStudy (allank):

I'd say no, because graphing that function gives the y intercept to be at y=0

OpenStudy (anonymous):

Can you explain further how you got that the y intercept is 0?

OpenStudy (anonymous):

Input x = 0?

OpenStudy (allank):

Not quite. Replace the f(x) in the function with y. y = x^3 - 33x^2 +216x - 2 Then enter that into your graphing calc

OpenStudy (accessdenied):

The y-intercept is where the graph crosses the y-axis, which means \(x=0\). Notice what happens to every variable term as \(x=0\), and how the constant is not affected.

OpenStudy (anonymous):

Yeah so to find the y intercept you simply do f(0)?

OpenStudy (accessdenied):

Yes. The actual graph of the function is incredibly steep at x=0, so the graph is kind of hard to read there. It isn't at y=0 though.

OpenStudy (allank):

True, @AccessDenied , had used an unsuitable scale.

OpenStudy (anonymous):

allank I think your mistake was that you looked at a graph on a calculator rather than solving for the y intercept

OpenStudy (allank):

@ConfusedAboutCalculus Graphing does work, though

OpenStudy (anonymous):

Didn't work for you this time :P

OpenStudy (allank):

Heheh, didn't use a good scale, and saw the graph as if crossing the y axis at zero. Gimme a break :)

OpenStudy (accessdenied):

Yeah, the slope around x=0 is in the hundreds for rise/run (285/1 at x=-1, 216/1 at x=0, 153/1 at x=1...). :P

OpenStudy (allank):

;)